OpenStudy (anonymous):

simultaneous equations with 3 variables. I kinda know what to do but I keep getting them wrong, so maybe I don't. Okay, 3 problems. 3x + y + 2z = 5 x - y - z = 0 2x + 3y = 2 What is x, y, z?

OpenStudy (anonymous):

x=1,y=0,z=1

OpenStudy (anonymous):

we can use gauss elimination method to solve it

OpenStudy (anonymous):

add the first 2 equations=> 4x+z=5 now multiply the second by 3 and add it to the 3rd. 5x-3z=2 now solve for x and z:z=5-4x =>5x-3(5-4x)=2 17x=17=>x=1 z=1 substitute in any of the equations to find y: 3+y+2=5 =>y=0 (x;y;z)=(1;0;1)

OpenStudy (anonymous):

okay just looking through your answer... thanks...

OpenStudy (anonymous):

okay cool hmmm I must be making little errors because my method seems ok... ah there are so many little calculations... well thankyou :)

OpenStudy (anonymous):

this is how i do it: choose any 2 of the equations to eliminate one variable choose other 2 equations to eliminate the same variable you end up with 2 equations in 2 unknowns,which you can solve easily. you can then find the eliminated variable from any of the 3 equations

OpenStudy (anonymous):

Or you can use cramer's rule.

OpenStudy (anonymous):

Cramers rule?

OpenStudy (anonymous):

it involves matrices and determinants to solve a system of equations: http://www.occc.edu/maustin/cramers_rule/cramer%27s%20rule.htm

OpenStudy (anonymous):

Oh okay I'll look forward to it when I get to matices :) Thanks

OpenStudy (anonymous):

haha oh dear it looks complex...

OpenStudy (anonymous):

Well, if you haven't been introduced to linear algebra yet, then I think solving systems by elimination and substitution will suffice.